# Estimation of Optimal Causal Bounds via Covariate-Assisted Optimal Transport

### v1

@@ 2.3 Conditional Optimal Transport
$\mathds{W}_{\textup{bc}}(\mu,\nu)=\min_{\pi\in\Pi(\mu_{Z},\nu_{Z})}\int\left\|{z-z^{\prime}}\right\|_{{2}}+\mathds{W}_{1}(\mu_{Y}^{z},\nu_{Y}^{z^{\prime}})~{}{\rm d}\pi(z,z^{\prime}).$
@@ 3 Basic Properties of COT Value
We first state a set of technical assumptions to enable the analysis below.
Assumption 3.1 (Compact covariate domain).
Assume that $\mathcal{Z}\subseteq[0,1]^{d_{Z}}$.
Assumption 3.2 (Moments of outcome).
Assume that, there is a constant $M\in[0,\infty)$, such that
$\sup_{z\in\mathcal{Z}}\mathbb{E}_{\mu}\left[\left\|{Y(0)}\right\|_{{2}}^{3}+\left\|{Y(1)}\right\|_{{2}}^{3}|Z=z\right]\leq M.$
Assumption 3.3 (Smoothness and bound of objective).
Assume that $h$ is $L_{h}$-Lipschitz continuous:
$\left|h(y_{0},y_{1})-h(y_{0}^{\prime},y_{1}^{\prime})\right|\leq L_{h}\left(\left\|{y_{0}-y_{0}^{\prime}}\right\|_{{2}}+\left\|{y_{1}-y_{1}^{\prime}}\right\|_{{2}}\right)\quad\forall(y_{0},y_{0}^{\prime},y_{1},y_{1}^{\prime})\in\mathcal{Y}^{4}.$
Assumption 3.4 (Lipschitz kernel).
Assume that there is a constant $L_{Z}>0$ such that for $k=0,1$,
$\mathds{W}_{1}\left(\mu_{Y(k)}^{z_{0}},\mu_{Y(k)}^{z_{1}}\right)\leq L_{Z}\left\|{z_{0}-z_{1}}\right\|_{{2}}\quad\textup{for}~{}\textup{$\mu_{Z}$-a.e.}~{}z_{0},z_{1}\in\mathcal{Z}.$
Assumption 3.1 and 3.2 impose technical tail conditions on the distribution $\mu_{Z}$ and $\mu_{Y(0)}^{z},\mu_{Y(1)}^{z}$. Specifically, we assume that the covariate $Z$ is supported on a compact subset of $\mathbb{R}^{d_{Z}}$, which is often satisfied in practice—for example, when $Z$ includes bounded physical measurements (e.g., age, height, or weight), socioeconomic indicators (e.g., income capped by a known threshold), or categorical variables embedded in Euclidean space. Assumption 3.3 and 3.4 are introduced to establish finite-sample convergence rates, and are aligned with standard regularity conditions in nonparametric estimation theory (e.g., [wasserman2006all, Theorem 5.65]).
@@ 3.1 Optimal Covariate-assisted PI Sets
In this section, we show that if we only have access to i.i.d. samples drawn from $\mu_{Y(0),Z}$ and $\mu_{Y(1),Z}$, respectively and no additional side information, then the interval $[V_{\textup{c}},\tilde{V}_{\textup{c}}]$ defined in (4) is the optimal PI set for $V^{*}$ that can be identified.
Theorem 3.1 (Optimal PI bounds).
Suppose that $h$ is a bounded continuous function, and the sample $((Y_{i}(0),Z_{i}(0)),i\in[n])$ i.i.d. follows $\mu_{Y(0),Z}$, $((Y_{j}(1),Z_{j}(1)),j\in[m])$ i.i.d. follows $\mu_{Y(1),Z}$. Then, (i) $V_{\textup{c}},\tilde{V}_{\textup{c}}$ are identifiable from the given samples, (ii) $[V_{\textup{c}},\tilde{V}_{\textup{c}}]$ is the smallest interval that contains all possible values of $\mathbb{E}_{\pi}[h(Y(0),Y(1))]$, where $\pi$ is aligned with distributions of the observed samples, that is, $\pi_{Y(0),Z}=\mu_{Y(0),Z},\pi_{Y(1),Z}=\mu_{Y(1),Z}$.
@@ 3.2 Stronger Topology for COT Functional
As indicated by Proposition 2.2, the estimation of $V_{\textup{c}}$ requires estimating the conditional distributions $\mu_{Y(0)}^{z},\mu_{Y(1)}^{z}$. Thus, a direct plug-in estimator does not yield consistent estimates of $V_{\textup{c}}$, a fact also observed in [chemseddine2024conditional]. To further illustrate this issue, we present an example adapted from [chemseddine2024conditional, Example 9].
Example 3.1 (Inconsistent estimator).
Let $m=n$ and $h=(y_{0}-y_{1})^{2}$. Suppose that under $\mu$, $Y(0)$, $Y(1)$, $Z$ are independent, and $\mu_{Y(0),Z}=\mu_{Y(1),Z}$, thus $V_{\textup{c}}=0$. Let $\hat{\mu}_{k}=\frac{1}{n}\sum_{i=1}^{n}\delta_{Y_{i}(k),Z_{i}(k)},k=0,1$. Further, assume that (i) the $i$-th covariate in the treated group is identical to that in the control group, i.e., $Z_{i}(1)=Z_{i}(0)~{}\forall i$, and (ii) the distribution $\mu_{Z}$ has a density. Then,
[…]
where $\hat{\Pi}_{\textup{c},n}=\left\{\pi:\pi_{Y(0),Z}=\hat{\mu}_{0},\pi_{Y(1),Z}=\hat{\mu}_{1}\right\}$. This is because we aim for a perfect match in $Z$ while almost surely the values of covariates in the treated group are distinct, i.e., $Z_{i}(1)\neq Z_{j}(1),i\neq j$. Therefore, $\hat{\Pi}_{\textup{c},n}$ only contains a single coupling: $\frac{1}{n}\sum_{i}\delta_{Y_{i}(0),Y_{i}(1),Z_{i}(0)}$. Thus, the objective simplifies to $\frac{1}{n}\sum_{i}(Y_{i}(1)-Y_{i}(0))^{2}$, which converges to the population quantity $\mathbb{E}_{\mu}\left[(Y(1)-Y(0))^{2}\right]$.
From a topological perspective, if $V_{\textup{c}}$ is viewed as a functional of $\mu_{Y(0),Z}$ and $\mu_{Y(1),Z}$, Example 3.1 implies that $V_{\textup{c}}$ is discontinuous under the weak topology. However, under mild assumptions, $V_{\textup{c}}$ becomes continuous at $\mu$ under the stronger topology induced by the adapted Wasserstein distance.
Theorem 3.2 (Continuity under $\mathds{W}_{\textup{bc}}$).
Let $(\nu_{n}=\nu_{Y(0),Y(1),Z,n},n\geq 1)$ be a sequence in $\mathcal{P}(\mathcal{Y}^{2}\times\mathcal{Z})$. Under Assumption 3.1 - 3.4, if  $\lim_{n\rightarrow\infty}\mathds{W}_{\textup{bc}}(\nu_{Y(k),Z,n},\mu_{Y(k),Z})=0$ for $k=0,1$, then
$\lim_{n\rightarrow\infty}\int\mathds{W}_{h}\left(\nu_{Y(0),n}^{z},\nu_{Y(1),n}^{z}\right){\rm d}\nu_{Z,n}(z)=\int\mathds{W}_{h}\left(\mu_{Y(0)}^{z},\mu_{Y(1)}^{z}\right){\rm d}\mu_{Z}(z)=V_{\textup{c}}.$
Remark 3.1 (Continuous kernel).
Assp. 3.4 can be relaxed to: $g_{k}(z)=\mu_{Y(k)}^{z}\ k=0,1$ are continuous under the weak topology. Since $\mathcal{Z}$ is compact, we get $g_{k}$ are uniformly continuous, and thus for any $\delta>0$, there is a constant $C(\delta)$ such that $\mathds{W}_{1}\left(\mu_{Y(k)}^{z_{0}},\mu_{Y(k)}^{z_{1}}\right)\leq\delta+C(\delta)\|z_{0}-z_{1}\|_{2}$, which is sufficient to show the result above.
In the next section, we construct an adapted empirical distribution that converges under $\mathds{W}_{\textup{bc}}$, which serves as the foundation for our adapted COT value estimator.
@@ 4.1.1 Adapted COT value estimator
In this section, we introduce our adapted COT value estimator in the setting where the treatment is independent of both the observed outcome and covariates. This setting is related to a scenario referred to as the Bernoulli design in causal inference, where $W_{i}$ are i.i.d. Bernoulli with a constant probability parameter. In this setting, the sample from the control group, $((Y_{i}(0),Z_{i}(0)),i\in[n])$, can be viewed as an i.i.d. sample drawn from $\mu_{Y(0),Z}$, with an analogous interpretation for the treated group.
Assumption 4.1 (Bernoulli design).
Assume that $((Y_{i}(0),Z_{i}(0)),i\in[n])$ are drawn i.i.d. from $\mu_{Y(0),Z}$, and $((Y_{j}(1),Z_{j}(1)),j\in[m])$ are drawn i.i.d. from $\mu_{Y(1),Z}$.
Inspired by [backhoff2022estimating], we introduce the adapted empirical distribution for constructing our COT value estimator. First, we introduce the map below for discretization.
Definition 4.1 (Cell-center projection).
Let $r>0$. For $n\geq 1$, partition the cube $[0,1]^{d_{Z}}$ into a disjoint union of $n^{rd_{Z}}$ smaller cubes, each with edge length $n^{-r}$. Define $\varphi_{r}^{n}:[0,1]^{d_{Z}}\to[0,1]^{d_{Z}}$ as the mapping that assigns each point to the center of the small cube in which it resides.
Definition 4.2 (Adapted empirical distribution).
@@
$\hat{\bm{\mu}}_{Y(0),Z(0),n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{Y_{i}(0),\varphi_{r}^{n}(Z_{i}(0))},\quad\hat{\bm{\mu}}_{Y(1),Z(1),m}=\frac{1}{m}\sum_{j=1}^{m}\delta_{Y_{j}(1),\varphi_{r}^{m}(Z_{j}(1))},$ (6)
where $\varphi_{r}^{n},\varphi_{r}^{m}$ are defined in Definition 4.1.
Compared to [backhoff2022estimating], our construction of $\hat{\bm{\mu}}$ has a key distinction: we discretize only the covariate $Z$, not the outcome $Y$, since discretization is required only for the variables being conditioned on.
If $z\mapsto\mu_{Y(i)}^{z}$ is continuous under the weak topology, the adapted empirical distributions (6) provide consistent approximations to the conditional distributions $\mu_{Y(k)}^{z},k=0,1$, thereby mitigating the issue that each sample of $Z(0)$ or $Z(1)$ is associated with only a single observed outcome $Y$ when $\mu_{Z}$ has a density. Further, [backhoff2022estimating] shows that $(\hat{\bm{\mu}}_{Y(k),Z(k),n},k=0,1)$ converges to their weak limits under $\mathds{W}_{\textup{bc}}$, which is aligned with the conditions introduced in Theorem 3.2.
Note that larger discretization cells contain more observations, and the resulting local distribution estimator is associated with smaller variance but larger bias. We can choose the cell size to balance the bias-variance trade-off, where the optimal cell size depends on factors including the sample size. As the sample size may differ between treatment and control groups, for example, when units are assigned to treatment with relatively low probability, we discretize each group separately to allow for distinct optimal cell sizes. In addition, in typical causal inference datasets, the covariate data for the treatment and control groups are not perfectly matched, also leading to a potential mismatch in the supports of the adapted empirical distributions in (6). To address the mismatch issue, we match discretized covariates $Z(0)$ from the control group with discretized covariates $Z(1)$ from the treated group that may differ slightly in value. This matching is formalized through a coupling with marginals $\hat{\bm{\mu}}_{Z(0),n}$ and $\hat{\bm{\mu}}_{Z(1),m}$. Based on this coupling, we define our adapted COT value estimator as follows.
Definition 4.3 (Adapted COT value estimator).
Given a coupling $\hat{\pi}_{n,m}\in\Pi(\hat{\bm{\mu}}_{Z(0),n},\hat{\bm{\mu}}_{Z(1),m})$, the COT value estimator is defined as
$\hat{V}_{n,m}\triangleq\int\mathds{W}_{h}(\hat{\bm{\mu}}^{z_{0}}_{Y(0),n},\hat{\bm{\mu}}^{z_{1}}_{Y(1),m}){\rm d}\hat{\pi}_{n,m}(z_{0},z_{1}).$ (7)
The COT value estimator achieves consistency as the sample size grows, provided that the coupling $\hat{\pi}_{n,m}$ converges to a joint distribution that exactly matches the limiting marginals.
Theorem 4.1 (Consistency).
Under Assp. 3.1 -  3.4, and assume that as $n,m\rightarrow\infty$, $\hat{\pi}_{n,m}$ weakly converges to $(\mathrm{id},\mathrm{id})_{\#}\mu_{Z}$ almost surely, then $\hat{V}_{n,m}$ converges to $V_{\textup{c}}$ almost surely.
@@ 4.1.2 Convergence Rate Analysis
Next, we provide the finite-sample convergence rate for the adapted COT value estimator. We first introduce an assumption to quantify the discrepancy between the estimated coupling $\hat{\pi}_{n,m}$ and the identity coupling $(\mathrm{id},\mathrm{id})_{\#}\mu_{Z}$, and to ensure that this error aligns with the granularity of the discretization, as controlled by the parameter $r$ in the map $\varphi_{r}^{n}$.
Assumption 4.2 (Coupling gap).
Assume that $\mathbb{E}[\int\left\|{z_{0}-z_{1}}\right\|_{{2}}{\rm d}\hat{\pi}_{n,m}(z_{0},z_{1})]\leq C_{\hat{\pi}}(n\wedge m)^{-r}$ for $n,m\geq 1$, where $C_{\hat{\pi}}$ is a constant independent of $n,m$.
Theorem 4.2 (Finite-sample complexity).
Under Assp. 3.14.1, and Assp. 4.2 with $r=\frac{1}{d_{Z}+2\vee d_{Y}}$, we have
$\mathbb{E}\left[|\hat{V}_{n,m}-V_{\textup{c}}|\right]\leq\bar{C}\gamma_{d_{Y},d_{Z}}(m\wedge n),$
@@
Here, $\bar{C}$ is a constant that depends on $d_{Z},d_{Y}$, $M$ (Assp. 3.2), $L_{h}$ (Assp. 3.3), and $L_{Z}$ (Assp. 3.4).
One natural choice of $\hat{\pi}_{n,m}$ that satisfies Assp. 4.2 is an optimal coupling of $\hat{\bm{\mu}}_{Z(0),n}$ and $\hat{\bm{\mu}}_{Z(1),m}$:
$\hat{\pi}_{n,m}\in\mathop{\mathrm{argmin}}_{\pi\in\Pi(\hat{\bm{\mu}}_{Z(0),n},\hat{\bm{\mu}}_{Z(1),m})}\mathbb{E}_{\pi}[\|Z(0)-Z(1)\|_{2}].$ (8)
Remark 4.1 (Robustness).
An additional advantage of using the coupling $\hat{\pi}_{n,m}$ defined above is its robustness to potential data perturbations that cause a mismatch between the covariate distributions of the treatment and control groups at the population level. Specifically, consider a perturbation such that $\mu_{Z}=\mu_{Z(0)}$, while the distribution of covariates in the treatment group satisfies $\mathds{W}_{1}(\mu_{Z},\mu_{Z(1)})\leq\epsilon$. Then, the adapted COT value estimator based on $\hat{\pi}_{n,m}$ enjoys a robustness guarantee in the sense that
$\mathbb{E}\left[|\hat{V}_{n,m}-V_{\textup{c}}|\right]\leq L_{h}L_{Z}\epsilon+\bar{C}\gamma_{d_{Y},d_{Z}}(m\wedge n).$
This advantage is further demonstrated empirically in Section 5.
Remark 4.2.
In the literature of COT, a variety of algorithmic strategies have been explored to deal with the inherent computational complexity of COT, including adversarial optimization via GANs [baptista2024representation], entropic regularization [baptista2024conditional], relaxations based on KL divergence [tabak2021data, manupriya2024consistent], and optimal transport approximations [chemseddine2024conditional]. While computational methods for COT have advanced significantly, rigorous statistical guarantees for value estimation have only begun to emerge in recent work [ji2023model, manupriya2024consistent, lin2025tightening]. Our work contributes to the statistical aspect of the COT literature by proposing a finite-sample estimator of the COT value with a provable convergence rate.
4.2 Covariate-Dependent Treatment Assignment

### v2

@@ 2.4 Adapted Wasserstein Distance and Its Topology
$\mathds{W}_{\textup{a}}(\mu,\nu)$
$=$ $\min_{\pi\in\Pi(\mu_{Z},\nu_{Z})}\int\left\|{z-z^{\prime}}\right\|_{{2}}+\mathds{W}_{1}(\mu_{Y}^{z},\nu_{Y}^{z^{\prime}})~{\rm d}\pi(z,z^{\prime}).$
Compared to the conventional Wasserstein distance, which measures proximity between full joint distributions of $(Y,Z)$ and $(Y^{\prime},Z^{\prime})$, the adapted Wasserstein distance quantifies discrepancies at the level of transition kernels. It therefore captures the transportation cost of conditional outcome laws while optimally aligning covariate marginals. As we show in Theorem 2, under mild regularity conditions, $V_{\textup{c}}$ becomes continuous with respect to this stronger, causally structured topology.
@@ 3 Basic Properties of COT Value
The following assumption enables our analysis.
Assumption 4 (Compact covariate domain).
Assume that $\mathcal{Z}\subseteq[0,1]^{d_{Z}}$, and $\mathcal{Y}\subseteq\mathbb{R}^{d_{Y}}$ is compact.
Assumption 5 (Continuous objective).
Assume that $h:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$ is a continuous function.
Here, we assume that the outcome $Y$ and the covariate $Z$ are supported on a compact set, which is often satisfied in practice—for example, when the outcome includes test scores, earnings, or health indicators and the covariate includes bounded physical measurements like age, height, or weight.
@@ 3.1 Optimal Covariate-Assisted PI Sets
In this section, we show that if we only have access to i.i.d. samples drawn from $\mu_{Y(0),Z}$ and $\mu_{Y(1),Z}$, respectively and no additional side information, then the interval $V_{\textup{c}},\tilde{V}_{\textup{c}}]$ defined in (3) is the optimal PI set for $V^{*}$ that can be identified. Note that, in the remaining parts of this work, we will use $((Y_{i}(w),Z_{i}(w))$ to denote the outcome and covariate values collected from the sample $i$ within the control group ($w=0$) and treatment group ($w=1$). Note that in this work, $Z_{i}(w)$ only contains pre-treatment covariates, and the $w$ in $Z_{i}(w)$ is the group index of the observed pre-treatment covariates.
Theorem 1 (Optimal PI bounds).
Under Assp. 4-5, suppose that the sample $((Y_{i}(w),Z_{i}(w)),i\in[n])$ i.i.d. follows $\mu_{Y(w),Z}$ for $w=0,1$. Then, we have (i) $V_{\textup{c}},\tilde{V}_{\textup{c}}$ are identifiable from the given sample, (ii) $V_{\textup{c}},\tilde{V}_{\textup{c}}]$ is the interval that exactly contains all possible values of $\mathbb{E}_{\pi}[h(Y(0),Y(1))]$, where $\pi$ satisfies: $\pi_{Y(0),Z}=\mu_{Y(0),Z},\pi_{Y(1),Z}=\mu_{Y(1),Z}$.
@@ 3.2 Stronger Topology for COT Functional
As indicated by Proposition 2, the estimation of $V_{\textup{c}}$ requires estimating the conditional distributions $\mu_{Y(0)}^{z},\mu_{Y(1)}^{z}$. Thus, a direct plug-in estimator does not yield consistent estimates of $V_{\textup{c}}$, a fact also observed in [16]. To further illustrate this issue, we consider an idealized case where the $i$-th covariate in the treated group is identical to that in the control group, i.e., $Z_{i}(1)=Z_{i}(0)~\forall i$ (e.g. a twin study), presenting an example adapted from [16, Example 9].
Example 2 (Inconsistent estimator).
Let $n_{0}=n_{1}$ and $h=|y_{0}-y_{1}|$. Suppose that under $\mu$, $Y(0)$, $Y(1)$, $Z$ are independent, and $\mu_{Y(0),Z}=\mu_{Y(1),Z}$, thus $V_{\textup{c}}=0$. Let $\hat{\mu}_{w}=\frac{1}{n_{w}}\sum_{i=1}^{n_{w}}\delta_{Y_{i}(w),Z_{i}(w)},w=0,1$. Further, assume that (i) $Z_{i}(1)=Z_{i}(0)~\forall i$, and (ii) the distribution $\mu_{Z}$ has a density. Then,
$\lim_{n_{0},n_{1}\rightarrow\infty}\min_{\pi\in\hat{\Pi}_{\textup{c}}}\mathbb{E}_{\pi}[h(Y(0),Y(1))]$
$=$ $\mathbb{E}_{\mu}[|Y(0)-Y(1)|]>0=V_{\textup{c}}\qquad\text{a.s.},$
where $\hat{\Pi}_{\textup{c}}=\left\{\pi:\pi_{Y(0),Z}=\hat{\mu}_{0},\pi_{Y(1),Z}=\hat{\mu}_{1}\right\}$. This is because we aim for a perfect match in $Z$ while almost surely the values of covariates in the treated group are distinct, i.e., $Z_{i}(1)\neq Z_{j}(1)$ for $i\neq j$. Therefore, $\hat{\Pi}_{\textup{c}}$ only contains a single coupling: $\frac{1}{n_{0}}\sum_{i}\delta_{Y_{i}(0),Y_{i}(1),Z_{i}(0)}$. Thus, the objective simplifies to $\frac{1}{n_{0}}\sum_{i}|Y_{i}(1)-Y_{i}(0)|$, which converges to the population quantity $\mathbb{E}_{\mu}\left[|Y(1)-Y(0)|\right]$.
From a topological perspective, if $V_{\textup{c}}$ is viewed as a functional of $\mu_{Y(0),Z}$ and $\mu_{Y(1),Z}$, Example 2 implies that $V_{\textup{c}}$ is discontinuous under the weak topology. However, under mild assumptions, $V_{\textup{c}}$ becomes continuous at $\mu$ under the stronger topology induced by the adapted Wasserstein distance (see Definition 1).
Theorem 2 (Continuity under $\mathds{W}_{\textup{a}}$).
Let $(\nu_{n}=\nu_{Y(0),Y(1),Z,n},n\geq 1)$ be a sequence in $\mathcal{P}(\mathcal{Y}^{2}\times\mathcal{Z})$. Suppose $g_{w}(z)=\mu_{Y(w)}^{z}\ w=0,1$ are continuous under the weak topology and $h$ is a continuous function. Under Assp. 4-5, if $\lim_{n\rightarrow\infty}\mathds{W}_{\textup{a}}(\nu_{Y(w),Z,n},\mu_{Y(w),Z})=0$ for $w=0,1$, then
$\lim_{n\rightarrow\infty}\int\mathds{W}_{h}\left(\nu_{Y(0),n}^{z},\nu_{Y(1),n}^{z}\right){\rm d}\nu_{Z,n}(z)$
$=$ $\int\mathds{W}_{h}\left(\mu_{Y(0)}^{z},\mu_{Y(1)}^{z}\right){\rm d}\mu_{Z}(z)=V_{\textup{c}}.$
In the next section, we construct an adapted empirical distribution that converges under $\mathds{W}_{\textup{a}}$, which serves as the foundation for our adapted COT value estimator.
@@ 4.1.1 Adapted COT Value Estimator
In this section, we introduce our adapted COT value estimator in the setting where the treatment is independent of both the observed outcome and covariates. This setting is related to a scenario referred to as the Bernoulli design (and also the randomized controlled trial) in causal inference, where $W_{i}$ are i.i.d. Bernoulli with a constant probability parameter. In this setting, the sample from the control group, $((Y_{i}(0),Z_{i}(0)),i\in[n_{0}])$, can be viewed as an i.i.d. sample drawn from $\mu_{Y(0),Z}$, with an analogous interpretation for the treated group.
Assumption 6 (Bernoulli design).
Assume that $((Y_{i}(w),Z_{i}(w)),i\in[n_{w}])$ are drawn i.i.d. from $\mu_{Y(w),Z}$ for $w=0,1$.
Inspired by [3], we introduce the adapted empirical distribution for constructing our COT value estimator. First, we introduce the map below for discretization.
Definition 2 (Cell-center projection).
Let $r>0$. For $n\geq 1$, partition the cube $0,1]^{d_{Z}}$ into a disjoint union of $n^{rd_{Z}}$ smaller cubes, each with edge length $n^{-r}$. Define $\varphi_{r}^{n}:[0,1]^{d_{Z}}\to[0,1]^{d_{Z}}$ as the mapping that assigns each point to the center of the small cube it resides in.
Definition 3 (Adapted empirical distribution).
@@
$\hat{\bm{\mu}}_{Y(w),Z(w),n_{w}}$ $=\frac{1}{n_{w}}\sum_{i=1}^{n_{w}}\delta_{Y_{i}(w),\varphi_{r}^{n_{w}}(Z_{i}(w))}$ (5)
for $w=0,1$, where $\varphi_{r}^{n}$ is defined in Definition 2.
Compared to [3], our construction of $\hat{\bm{\mu}}$ has a key distinction: we discretize only the covariate $Z$, not the outcome $Y$, since discretization is required only for the variables being conditioned on.
If $z\mapsto\mu_{Y(w)}^{z}$ is continuous under the weak topology, the adapted empirical distributions (5) provide consistent approximations to the conditional distributions $\mu_{Y(w)}^{z},w=0,1$, thereby mitigating the issue that each sample of $Z(0)$ or $Z(1)$ is associated with only a single observed outcome $Y$ when $\mu_{Z}$ has a density. Further, [3] suggests that $(\hat{\bm{\mu}}_{Y(w),Z(w),n_{w}},w=0,1)$ converges to their weak limits under $\mathds{W}_{\textup{a}}$, which is aligned with the conditions introduced in Theorem 2.
Note that larger discretization cells contain more observations, and the resulting local distribution estimator is associated with smaller variance but larger bias. We can choose the cell size to balance the bias-variance trade-off, where the optimal cell size depends on factors including the sample size. As the sample size may differ between treatment and control groups, for example, when units are assigned to treatment with relatively low probability, we discretize each group separately to allow for distinct optimal cell sizes. In addition, in typical causal inference datasets, the covariate data for the treatment and control groups are not perfectly matched, also leading to a potential mismatch in the supports of the adapted empirical distributions in (5). To address this issue, we match the discretized covariates $Z(0)$ and $Z(1)$ that may differ in value. This matching is formalized as a coupling with marginals $\hat{\bm{\mu}}_{Z(0),n_{0}}$ and $\hat{\bm{\mu}}_{Z(1),n_{1}}$. Based on this coupling, we define our adapted COT value estimator as follows.
Definition 4 (Adapted COT value estimator).
Given a coupling $\hat{\pi}_{n_{0},n_{1}}\in\Pi(\hat{\bm{\mu}}_{Z(0),n_{0}},\hat{\bm{\mu}}_{Z(1),n_{1}})$, the COT value estimator is defined as
$\hat{V}_{n_{0},n_{1}}$ $\triangleq\int\mathds{W}_{h}(\hat{\bm{\mu}}^{z_{0}}_{Y(0),n_{0}},\hat{\bm{\mu}}^{z_{1}}_{Y(1),n_{1}}){\rm d}\hat{\pi}_{n_{0},n_{1}}(z_{0},z_{1}).$
The COT value estimator achieves consistency as the sample size grows, provided that the coupling $\hat{\pi}_{n_{0},n_{1}}$ converges to a joint distribution that exactly matches the limiting marginals.
Theorem 3 (Consistency).
Under Assp. 4-6, if as $n_{0},n_{1}\rightarrow\infty$, $\hat{\pi}_{n_{0},n_{1}}$ weakly converges to $(\mathrm{id},\mathrm{id})_{\#}\mu_{Z}$ almost surely, then $\hat{V}_{n_{0},n_{1}}$ converges to $V_{\textup{c}}$ almost surely.
@@ 4.1.2 Convergence Rate Analysis
Next, we provide the finite-sample convergence rate for the adapted COT value estimator. We introduce the following assumptions.
Assumption 7 (Smooth objective).
Assume that $h$ is $L_{h}$-Lipschitz continuous: for $\forall y_{0},y_{0}^{\prime},y_{1},y_{1}^{\prime}\in\mathcal{Y}$,
$\left|h(y_{0},y_{1})-h(y_{0}^{\prime},y_{1}^{\prime})\right|\leq L_{h}\left(\left\|{y_{0}-y_{0}^{\prime}}\right\|_{{2}}+\left\|{y_{1}-y_{1}^{\prime}}\right\|_{{2}}\right).$
Assumption 8 (Lipschitz kernel).
Assume that there is a constant $L_{Z}>0$ such that for $\textup{$\mu_{Z}$-a.e.}~z_{0},z_{1}\in\mathcal{Z}$,
$\mathds{W}_{1}\left(\mu_{Y(w)}^{z_{0}},\mu_{Y(w)}^{z_{1}}\right)\leq L_{Z}\left\|{z_{0}-z_{1}}\right\|_{{2}}\ \ w=0,1.$
Assp. 7-8 are aligned with standard regularity conditions in nonparametric estimation theory (e.g., [51, Theorem 5.65]).
Assumption 9 (Coupling gap).
Assume that $\forall n_{0},n_{1}$ $\mathbb{E}[\int\left\|{z_{0}-z_{1}}\right\|_{{2}}{\rm d}\hat{\pi}_{n_{0},n_{1}}(z_{0},z_{1})]\leq C_{\hat{\pi}}(n_{0}\wedge n_{1})^{-r}$, where $C_{\hat{\pi}}$ is a constant independent of $n_{0},n_{1}$.
Assp. 9 enables quantifying the discrepancy between the estimated coupling $\hat{\pi}_{n_{0},n_{1}}$ and the identity coupling $(\mathrm{id},\mathrm{id})_{\#}\mu_{Z}$, and to ensure that this error aligns with the granularity of the discretization, as controlled by the parameter $r$ in the map $\varphi_{r}^{n}$.
Theorem 4 (Finite-sample complexity).
Under Assp. 4-9 with $r=\frac{1}{d_{Z}+2\vee d_{Y}}$, we have
$\mathbb{E}\left[|\hat{V}_{n_{0},n_{1}}-V_{\textup{c}}|\right]\leq\bar{C}\gamma_{d_{Y},d_{Z}}(n_{0}\wedge n_{1}),$
@@
Here, $\bar{C}$ is a constant that depends on $d_{Z},d_{Y}$, $L_{h}$ (Assp. 7), and $L_{Z}$ (Assp. 8).
One natural choice of $\hat{\pi}_{n_{0},n_{1}}$ that satisfies Assp. 9 is an optimal coupling of $\hat{\bm{\mu}}_{Z(0),n_{0}}$ and $\hat{\bm{\mu}}_{Z(1),n_{1}}$:
$\hat{\pi}_{n_{0},n_{1}}\in\mathop{\mathrm{argmin}}_{\pi\in\Pi(\hat{\bm{\mu}}_{Z(0),n_{0}},\hat{\bm{\mu}}_{Z(1),n_{1}})}\mathbb{E}_{\pi}[\|Z(0)-Z(1)\|_{2}].$ (6)
An additional advantage of the coupling $\hat{\pi}_{n_{0},n_{1}}$ (6) is that it enjoys a robustness guarantee against potential data perturbations to the covariate distributions.
Theorem 5 (Robustness).
Define $\hat{\pi}_{n_{0},n_{1}}$ by (6). Suppose that the sample of $Z(w)$ is i.i.d. drawn from $\mu^{\prime}_{w}$, satisfying $\mathds{W}_{1}(\mu_{Z},\mu^{\prime}_{w})\leq\epsilon$ for $w=0,1$. Then,
$\mathbb{E}\left[|\hat{V}_{n_{0},n_{1}}-V_{\textup{c}}|\right]\leq 2L_{h}L_{Z}\epsilon+\bar{C}\gamma_{d_{Y},d_{Z}}(n_{0}\wedge n_{1}).$
4.2 Covariate-Dependent Treatment